A theater group made appearances in two cities. The hotel charge before tax in the second city was $1500 higher than in the first. The tax in the first city was 6%,and the tax in the second city was 9%. The total hotel tax paid for the two cities was $735. How much was the hotel charge in each city before tax?
Answer by lwsshak3(11628)

A theater group made appearances in two cities. The hotel charge before tax in the second city was $1500 higher than in the first. The tax in the first city was 6%,and the tax in the second city was 9%. The total hotel tax paid for the two cities was $735. How much was the hotel charge in each city before tax?
***
let x=hotel charge before tax in first city
x+1500=hotel charge before tax in second city
6%x+9%(x+1500)=735
.06x+.09x+135=735
.15x=600
x=4000
x+1500=5500
hotel charge before tax in first city=$4000
hotel charge before tax in second city=$5500